OpenStudy (anonymous):

Identify ALL true statements. FORMEDALLLL y=mx+b is the point-slope form of a linear equation y=-4x^2+3x-1 is a quadratic equation ALL equations MUST have an equal sign ALL expressions MUST have an equal sign

OpenStudy (anonymous):

a is correct

OpenStudy (anonymous):

except the last , the rest are true

OpenStudy (anonymous):

the first statement is truth the second one use wrong; it's not a quadratic equation it's a parabola.

OpenStudy (anonymous):

and the two last are true too

OpenStudy (anonymous):

Lol now im confused you all have different answers haha

OpenStudy (anonymous):

A is trie for sure got that one.

OpenStudy (anonymous):

b is true

OpenStudy (anonymous):

it's not a quadratic equation a quadratic equation must be in the form \[ax^{2}+bx+c=0\]

OpenStudy (anonymous):

the first (x,y) is the point, m is the slope a quadratic equation is a equation in second degree of x. it doesnt matter that it is equal to y the third equations must have an eqial to sign expressions can have greater than, less than and all that also

OpenStudy (anonymous):

yeah u got a point

OpenStudy (anonymous):

and a parabola must be of the form y=4ax

OpenStudy (anonymous):

so which ones do I go for than ?

OpenStudy (anonymous):

i'll say all except the last are correct

OpenStudy (anonymous):

the second one and third one . Y=mx+b is slope intercept form and expressions don't require equal signs

OpenStudy (anonymous):

The "point-slope" form of the equation of a straight line is: y - y1 = m(x - x1)

OpenStudy (anonymous):

Alright thank you all so much.

OpenStudy (anonymous):

you can modified it a parabola is also in this form \[(y-y_{0})^{2}=2p(x-x_{0}) \] \[(x-x_{0})^{2}=2p(y-y_{0}) \]

OpenStudy (anonymous):

u r right! the graph of a quadratic equation is a parabola what i meant to say was that it doesnt matter that there is a y it is still a quadratic

OpenStudy (anonymous):

yeah :)